# SAM3D: Zero-Shot Semi-Automatic Segmentation in 3D Medical Images with the Segment Anything Model

### v1

@@ Abstract
We introduce SAM3D, a new approach to semi-automatic zero-shot segmentation of 3D images building on the existing Segment Anything Model. We achieve fast and accurate segmentations in 3D images with a four-step strategy comprising: volume slicing along non-orthogonal axes, efficient prompting in 3D, slice-wise inference using the pretrained SAM, and recoposition and refinement in 3D. We evaluated SAM3D performance qualitatively on an array of imaging modalities and anatomical structures and quantify performance for specific organs in body CT and tumors in brain MRI. By enabling users to create 3D segmentations of unseen data quickly and with dramatically reduced manual input, these methods have the potential to aid surgical planning and education, diagnostic imaging, and scientific research.
@@
Zero-shot segmentation 3D segmentation Semi-automatic segmentation
@@ 1 Introduction
Current automated approaches to image segmentation predominantly use deep learning models trained on vast quantities of labeled data. In medicine, these models commonly achieve high performance through hyperfixation: they train on a single anatomical region imaged using a single modality. There are numerous downsides to such an approach. Training a model first requires gathering a large amount of annotated data, which can be time consuming and expensive. And, because training datasets are imperfect, these models are also susceptible to brittleness and bias. When presented with images that deviate only slightly from those in the training set, often in ways unnoticeable by humans, the models can fail unexpectedly [23, 17]. Lastly, many medical imaging modalities acquire 3D images, which drastically increases the difficulty of storing, annotating, and processing sufficiently large and diverse datasets.
Because of this, there is a growing interest in semi-automatic approaches for general image segmentation. These methods offer a compelling compromise: by accepting a small decrease in speed and convenience compared to their fully-automated counterparts, we can have both increased reliability and greater generalizability. Previous proposed approaches to semi-automatic segmentation include early thresholding and region-growing methods for 2D and 3D, which rely on intensity values to delineate structures of interest [11]. Watershed and active contour methods segment on the basis of intensity gradients [1, 8, 30], and atlas and multi-atlas-based segmentation combines registration with prior knowledge in the form of an idealized reference image or set of images and labels [2, 16, 27]. With the advent of deep learning-based segmentation and the release of off-the-shelf models and pipelines [7, 28], high-quality automatic segmentation is easier than ever, but these models still require large amounts of domain-specific training data, so researchers have sought to combine the speed and performance of deep learning methods with the generalizability of previous semi-automatic algorithms [4, 24].
@@
Figure 1: An overview of the segmentation method comprising: (a) an initial choice of the number of axes to slice across (only one transform of four for cubic is shown here for clarity), (b) user prompting input as polylines to medial slices and the calculation of prompt intersection points, (c) inference on 2D slices with the pretrained SAM, (d) point extraction and user-mediated refinement of the dense point cloud, and (e) voxelization of the point cloud and user-mediated final mask refinement.
@@ 2 Methods
The method we propose for 3D segmentation is conceptually simple. Noting that SAM attains high zero-shot performance, we only need to slice a 3D image volume into a set of 2D images, add appropriate prompting to these images, segment using the pretrained model, and recompose the results into a 3D mask.
In practice, multiple decisions regarding the strategy of volume decomposition, prompting, and recomposition affect both the quality of the final 3D mask as well as the amount of human effort and time required. We select planar slices along a predefined set of n rotationally equispaced axes oriented with the vertices of a platonic solid (Figure 1a). Doing so ensures a near uniform sampling of the entire imaging region while also allowing the 2D SAM model to see each location in the imaging region n times, in essence giving it that many opportunities to segment the region correctly. Depending on the complexity of the anatomy being segmented, more or fewer axes may be called for, so the first user input is a choice of the number and distribution of axes: octahedral=3, cubic=4, icosahedral=6, and dodecahedral=10. (Once symmetries are removed, tetrahedral slicing is equivalent to cubic slicing, and is therefore omitted.)
Adding point prompts to every one of these slices, usually numbering in the hundreds, is infeasible, but it is also unnecessary. We choose only the median plane along each slicing axis to add prompting to. In contrast to SAM, which takes positive and negative prompts as 0D points, we prompt with 1D polylines that are fully contained within the structure of interest (positive) or fully outside of it (negative) (Figure 1b). We find that adding prompting to a 2D image with polylines is nearly as fast as prompting with points. But, by transforming these lines to 3D and calculating their intersections with all slicing planes along the other axes, we can generate point prompts over the entire 3D volume in an equivalent time as it would take to prompt 3-10 2D images, the number depending on the choice of transformation axes.
The resulting slices and prompt points are fed into the basic SAM model to predict a set of 2D masks suspended in 3D space (Figure 1c). As the segmentation predictions from SAM are not 100% accurate, a method for post-processing the aggregated results is needed. We convert a uniformly-sampled subset of voxels in each 2D mask to points in 3D, forming a dense 3D point cloud. While erroneous segmentation predictions in 2D result in errant points, these are extremely sparse when transformed to 3D, and can be easily and efficiently pruned with outlier detection filters (Figure 1d). The resulting dense point cloud can be voxelized and further refined using a combination of binary dilation, erosion, and hole-filling to obtain the final mask prediction (Figure 1e).
While refining the point clouds and voxel masks is both quick and intuitive, we found that the choice of parameters to use depends in part on the anatomical structure-of-interest, so each of these steps involves a small amount of user interaction to specify the aggressiveness of point cloud outlier removal and the degree of dilation and/or erosion, which affects smoothness. For a typical structure in a $256\times 256\times 256$ voxel image, this entire process takes on the order of minutes, roughly half of which is devoted to active user prompting and post-processing, the other half to image transformations and model inference.
@@
We further quantified the accuracy of these segmentations using two popular 3D medical imaging datasets: one is the Beyond the Cranial Vault (BTCV) dataset for organ segmentation in abdominal CT [13], the other is the Brain Tumor Segmentation (BraTS) dataset for glioblastoma in 4 MRI contrasts [22]. Both datasets have manually labeled masks, which serve as ground truth comparisons.
Figure 2: Visualizing diverse segmentation performance. (a) The pelvis and sacral spine in an abdominal pelvic CT. (b) Skeleton in an ex vivo CT scan of the hand. (c) Cervical vertebra 3 in a CT scan of the head and neck [15]. (d) Lungs in a chest CT scan [20]. (e) Lungs, liver, and kidneys in an abdominal CT scan [13]. (f) Oxygenated blood pool in a cardiac CT scan. (g) Glioblastoma tumor and edema regions in a FLAIR MRI scan of the brain [22]. (h) Lateral ventricles, cerebellum, and brain stem in a T1 MRI scan of the brain. (i) Left ventricular outflow tract, aortic valve, and aortic root in a 3D mid-esophageal transesophageal echocardiography scan of the heart. (j) A tumor lesion in 3D breast ultrasound [25]. (k) Axonol neurons in volumetric scanning electron microscopy (SEM) of a region of the hippocampus [18]. (l) Central vacuoule in volumetric SEM of a tobacco leaf cell [29].
Figure 3: Quantification of segmentation accuracy on benchmark datasets and additional experiments. (a) Dice score and intersection over union calculated for the lung and liver masks ($n=16$) on the BTCV dataset. (b) Dice score and intersection over union calculated for the tumor region (enhanced + nonenhanced tumor/necrotic) and the tumor+edema regions ($n=8$) for 4 MRI contrasts on the BraTS dataset. (c) Segmentation predictions for tumor and tumor+edema regions across four scans are highly consistent with each other, showing that predictions capture anatomical features and are robust to contrast, brightness, and textural changes. (d) An analysis of segmentation accuracy and time as a function of the number of transforms chosen for liver segmentation in the BTCV dataset ($n=3$) and tumor segmentation in the BraTS dataset ($n=3$) suggests that the optimal number of transforms to use depends heavily on the anatomical structure to segment. (e) Representative segmentation predictions (opaque) for a liver and brain tumor overlaid with the manually labeled ground-truth masks (translucent).
@@ 3 Results
Table 1: A comparison of segmentation times for 5 structures across 2 common datasets.
Dataset Segmentation Target Time Taken (Mean +/- Std.)
BTCV[13] Liver 4 min +/- 60 sec
Lungs 3 min 38 sec +/- 51 sec
Kidneys 3 min 50 sec +/- 1 min 47 sec
BraTS[22] ET + NCT/NET 4 min 14 sec +/- 1 min 5 sec
ET + NCT /NET+ ED 7 min 50 sec +/- 28 sec
@@ 4 Discussion
We introduce a new and efficient method for prompting and inference of a pretrained semi-automatic segmentation model and demonstrate strong results on a variety of 3D images. It is worthwhile to ask why SAM performs as well as it does in this task. Given that none of the images we evaluate on particularly resemble any images in the segment anything dataset, it is unreasonable to believe that the model is performing segmentations based on a semantic understanding of the data. Rather, we hypothesize that the SAM essentially mimics its much simpler precursor algorithms by relying on regional intensities, gradients, and textures to construct a hierarchical representation of the image which, when combined with prompting, can yield an accurate segmentation. While using a 2D model for 3D segmentation may seem unintuitive, this approach carries numerous benefits. It allows us to escape the memory and computational cost associated with processing 3D images. Most importantly, it allows us to leverage existing 2D datasets which, due to being easier to acquire and annotate, are far more comprehensive than their 3D counterparts.
We envision numerous uses of the proposed method. In scientific research, a method for fast semi-automatic 3D segmentation could be used in data-limited regimes or as a means to initially label training datasets. In medicine, 3D segmentation has applications in surgical planning, diagnostic imaging, and radiomics. The use of this tool could save physicians time while mitigating the risks of bias and unpredictability that plague fully automated models.
There are some notable limitations of SAM3D. As is the case for many segmentation models, including the base SAM, SAM3D has poor performance when labeling thin and branching structures. This is doubly the case in 3D, as prompt planes are sparse, so there are few opportunities to add prompting to these structures. A second limitation is that, because the base 2D segmentation model has not been trained on medical images, it lacks any relevant domain knowledge. This is potentially also an opportunity: multiple groups have taken the base SAM model and fine tuned it on a medical domain [19, 6, 21], and demonstrate improved results compared to the base model. Inserting these models as drop-in replacements for the SAM in our method could further improve segmentation performance and efficiency. Lastly, while we designed a rudimentary user interface for prompting and post-processing, broader use of these and similar tools would likely require a more polished interface. Other groups have shown success integrating SAM into existing medical visualization tools [14], and this is an approach we are considering.
@@ 5 Conclusion
SAM3D is a semi-automated, zero-shot 3D segmentation model capable of high performance across a range of structures and images. By enabling users to quickly, easily, and accurately segment 3D images, we hope that these methods will aid clinicians and researchers, accelerate the creation of large-scale 3D datasets, and spur development in general 3D segmentation models.

### v2

@@ Abstract
We introduce SAM3D, a new approach to semi-automatic zero-shot segmentation of 3D images building on the existing Segment Anything Model. We achieve fast and accurate segmentations in 3D images with a four-step strategy involving: user prompting with 3D polylines, volume slicing along multiple axes, slice-wide inference with a pretrained model, and recomposition and refinement in 3D. We evaluated SAM3D performance qualitatively on an array of imaging modalities and anatomical structures and quantify performance for specific structures in abdominal pelvic CT and brain MRI. Notably, our method achieves good performance with zero model training or finetuning, making it particularly useful for tasks with a scarcity of preexisting labeled data. By enabling users to create 3D segmentations of unseen data quickly and with dramatically reduced manual input, these methods have the potential to aid surgical planning and education, diagnostic imaging, and scientific research.
@@
Zero-shot segmentation, 3D segmentation, Semi-automatic segmentation
@@ 1 INTRODUCTION
Current automated approaches to image segmentation predominantly use deep learning models trained on vast quantities of labeled data. In medicine, these models typically achieve high performance through hyperfixation: they train on a single anatomical region imaged using a single modality. While narrowly effective, such an approach requires gathering a large amount of annotated data, which can be time consuming and expensive. And, because training datasets are imperfect, these models are also susceptible to brittleness and bias [6, 7]. Lastly, many medical imaging modalities acquire 3D images, which drastically increases the difficulty of storing, annotating, and processing sufficiently large and diverse datasets.
Because of this, there is a growing interest in semi-automatic approaches for general image segmentation. These methods offer a compelling compromise: by accepting a small decrease in speed and convenience compared to their fully-automated counterparts, we can have both increased reliability and greater generalizability. Early approaches to semi-automatic segmentation include thresholding and region-growing methods for 2D and 3D [8], watershed and active contour methods [9, 10, 11], and atlas and multi-atlas-based segmentation methods [12, 13, 14]. With the advent of deep learning-based segmentation and the release of off-the-shelf models and pipelines [15, 16], high-quality automatic segmentation is easier than ever. However, these models still require large amounts of domain-specific training data, so researchers have sought to combine the speed and performance of deep learning methods with the generalizability of previous semi-automatic algorithms [17, 18].
@@ 2 Methods
In principle, extending the zero-shot performance of SAM to 3D images is fairly simple. We can slice a 3D image into many 2D images, add appropriate prompting to these images, segment using the pretrained model, and recompose the results into a 3D mask. In practice, multiple decisions regarding the strategy of volume decomposition, prompting, and recomposition affect both the quality of the final 3D mask as well as the amount of human effort and time required.
When adding prompting to a 3D image, there is a trade off between precision and time; on one hand, prompting every 2D slice individually can yield accurate segmentations but it is very time consuming. On the other, using the same prompt for each 2D slice, such as defining a single 3D bounding box, saves time but introduces ambiguities when inferencing the model. A second major problem occurs in recomposition. Segmenting many individual slices multiplies the likelihood of a bad segmentation prediction, and these errors appear as unrealistic artifacts when images are recomposed in 3D. In the 2D case, these can be easily fixed with revised prompting, but this is not efficient for volumetric images.
Figure 1: An overview of the segmentation method comprising: polyline prompting on a 3D image, slicing along rotationally equispaced axes, 2D inference using SAM, recomposition into a dense point cloud, and voxelization/meshing.
We address these problems using a novel prompting strategy and a novel slicing strategy. User prompts take the form of polylines (positive and negative) added to slices of a 3D volume. At inference, we calculate the intersections between 1D line segments and the 2D plane we are segmenting to obtain a set of positive and negative prompt points which are then used to condition the pretrained model. In this manner, we can add precise prompting to an entire 3D volume consisting of hundreds of slices while manually prompting fewer than 10 slices.
To address errors that occur during segmentation, we resort to a multi-slicing strategy that produces redundancy in the segmentation outputs. While a typical pipeline for 3D segmentation with a 2D model might traverse along a single axis, we traverse along a predefined set of axes. The resulting planes uniformly cover the 3D volume and, because they intersect, they allow the pretrained segmentation model to see the same spatial location in the volume multiple times, giving it that many chances to segment it correctly. In order to ensure an unbiased distribution of slices, we select orthogonal planes along rotationally equispaced axes. Depending on the complexity of the anatomy being segmented, more or fewer axes may be called for, so the user has the option of using between 3 and 10 independent axes, resulting in a small trade off between accuracy and runtime. And, by taking larger steps along each axis, we can reduce the total number of segmentations the model needs to perform, further reducing runtime.
Recomposing 2D masks back into 3D occurs in two steps. First, we convert all segmentation masks into 2D point clouds and transform the points back into the global 3D reference frame. Second, we filter to remove outlier points in low-density regions. This step exploits the redundancy of the 2D planes: areas in erroneously segmented 2D slices are sparse in 3D space, allowing them to be efficiently removed without affecting the correctly segmented 3D shape. The final point cloud is voxelized and turned into an image mask or an object mesh.
For a typical structure in a $256^{3}$ voxel image, this entire process takes a few minutes, roughly a third of which is devoted to user prompting and post-processing and two thirds to image transformations and model inference. All experiments and evaluations were performed on a workstation with a single Nvidia 3090 gpu.
@@
We further quantified the accuracy of these segmentations using two 3D medical imaging datasets: the Beyond the Cranial Vault (BTCV) dataset for organ segmentation in abdominal CT [20], and the Brain Tumor Segmentation (BraTS2021) dataset for glioblastoma in 4 MRI contrasts [21]. Both datasets have manually labeled masks which serve as ground truth comparisons. In addition, we benchmarked the performance of our method against the popular MedSAM model [22] on a range of tasks including brain tumor and metastases segmentation in MRI, pelvic segmentation in CT, and aorta segmentation in transesophageal echocardiography (TEE).
Figure 2: Visualizing diverse segmentation performance. (a) Pelvis and sacral spine in CT. (b) Skeleton in ex vivo CT. (c) Cervical vertebra 3 in CT[23]. (d) Lungs in chest CT[24]. (e) Lungs, liver, and kidneys in abdominal CT[20]. (f) Oxygenated blood pool in CT angiogram. (g) Glioblastoma tumor and edema in FLAIR MRI[21]. (h) Lateral ventricles, cerebellum, and brain stem in T1 MRI. (i) Left ventricular outflow tract, aortic valve, and aortic root in 3D TEE. (j) Tumor lesion in 3D breast ultrasound[25]. (k) Hippocampal axonol neurons in volumetric scanning electron microscopy (SEM)[26]. (l) Tobacco leaf cell central vacuoule in volumetric SEM[27].
Figure 3: Quantification of segmentation accuracy on benchmark datasets, comparison against MedSAM, and additional experiments. (a) Dice score calculated for the lung and liver masks ($n=16$) on the BTCV dataset. (b) Dice score calculated for the tumor region (enhanced + nonenhanced tumor/necrotic) and the tumor+edema regions ($n=8$) for 4 MRI contrasts on the BraTS dataset. (c) Dice scores for segmentation performance on brain metastases ($n=4$), pelvis ($n=3$), and aorta ($n=4$) for both our method (SAM3D) and MedSAM. (d) An analysis of segmentation accuracy and time as a function of the number of transforms chosen for liver segmentation in the BTCV dataset ($n=3$) and tumor segmentation in the BraTS dataset ($n=3$) suggests that the optimal number of transforms to use depends heavily on the anatomical structure to segment. (e-g) Representative segmentation predictions shown in 2D and 3D in three zero-shot tasks for our method and MedSAM.
@@ 3 Results
We also compared the zero-shot performance of our model to that of the popular MedSAM model in 3 difficult segmentation tasks. To the best of our knowledge, MedSAM was not explicitly trained on either pelvic segmentation in CT or aortic segmentation in TEE despite having trained on a large variety of data including MRI, CT, and ultrasound. SAM, the model our method uses, was not trained on any medical images. Ground truth masks for each modality were created manually using itk-SNAP. We found that our model achieves significantly higher accuracy than MedSAM in the pelvic segmentation and aortic segmentation tasks (Figure 3c). In the brain tumor segmentation task, for which MedSAM was explicitly trained, the models performed comparably. Our method requires less time on average to perform a segmentation compared to MedSAM. Both models were orders of magnitude faster than manual segmentation, which took on average over an hour to produce a single 3D mask.
@@
Table 1: A comparison of segmentation times.
Method Dataset Segmentation Target Time Taken (Mean +/- Std.)
SAM3D BTCV[20] Liver 4 min +/- 60 sec
SAM3D BTCV[20] Lungs 3 min 38 sec +/- 51 sec
SAM3D BTCV[20] Kidneys 3 min 50 sec +/- 1 min 47 sec
SAM3D BraTS[21] ET + NCT/NET 4 min 14 sec +/- 1 min 5 sec
SAM3D BraTS[21] ET + NCT /NET+ ED 7 min 50 sec +/- 28 sec
SAM3D MRI Metastases 3 min 31 sec +/- 1 min 57 sec
SAM3D CT Pelvis 6 min 22 sec +/- 1 min 1 sec
SAM3D TEE Aorta 4 min 21 sec +/- 42 sec
MedSAM[22] MRI Metastases 7 min 40 sec +/- 1 min 17 sec
MedSAM[22] CT Pelvis 7 min 2 sec +/- 49 sec
MedSAM[22] TEE Aorta 7 min 3 sec +/- 20 sec
Manual CT Pelvis 1 hour min 28 min +/- 20 min
Manual TEE Aorta 1 hour 20 min +/- 44 min
@@ 4 Discussion
We introduce a new method for zero-shot semi-automatic 3D image segmentation that leverages a pretrained 2D segmentation model and demonstrate strong results on a variety of 3D images.
Despite SAM not having seen any 2D or 3D medical imaging data during training, it still performs highly, suggesting a few insights. First, it seems that a learning a true semantic understanding of the data might not be necessary; regional intensities, gradients, and textures, combined with sufficient prompting, are enough to reliably segment a wide variety of structures found in medical and scientific images. (This complements related work in synthetic data for image learning [28, 29].)
Second, while it may seem unintuitive to use a 2D model to perform 3D segmentation, the approach carries numerous benefits. 2D data is far less expensive to collect and store and far easier to train models on, and this means that the best 2D models, such as SAM, have seen orders of magnitude more data than the best 3D models. For zero-shot segmentation, and possibly many other tasks, the benefit of seeing more out-of-domain data may actually outweigh the benefit of seeing data that is closely aligned with the desired task.
There are some notable limitations of SAM3D. As is the case for many segmentation models, including the base SAM, SAM3D has poor performance when labeling thin and branching structures. A second limitation is that, because the base 2D segmentation model has not been trained on medical images, it lacks any relevant domain knowledge. Multiple groups have taken the base SAM model and fine-tuned it on a medical domain [22, 30, 31, 32], and demonstrate improved results compared to the base model. Inserting these models as drop-in replacements for the SAM in our method could further improve segmentation performance and efficiency.
@@ 5 Conclusion
SAM3D is a semi-automated, zero-shot 3D segmentation model capable of producing accurate segmentations across a range of structures and images. There are numerous potential uses for the proposed method. In scientific research, semi-automatic segmentation could be used in data-limited regimes or as a means to initially label training datasets. In medicine, 3D segmentation has widespread applications in surgical planning, diagnostic imaging, and radiomics. The use of this tool could save physicians time while mitigating the risks of bias and unpredictability that plague fully automated models. By enabling faster, easier, and more accurate 3D segmentation, we hope that these methods will aid clinicians and researchers, accelerate the creation of large-scale 3D datasets, and spur development in general 3D segmentation models.